# A Survey on Diffusion Language Models

### v1

@@ Abstract
Diffusion Language Models (DLMs) are rapidly emerging as a powerful and promising alternative to the dominant autoregressive (AR) paradigm. By generating tokens in parallel through an iterative denoising process, DLMs possess inherent advantages in reducing inference latency and capturing bidirectional context, thereby enabling fine-grained control over the generation process. While achieving a several-fold speed-up, recent advancements have allowed DLMs to show performance comparable to their autoregressive counterparts, making them a compelling choice for various natural language processing tasks. Despite their growing prevalence, DLMs present challenges and opportunities that warrant further exploration, requiring a detailed and systematic understanding of their principles, techniques, and limitations. In this survey, we provide a holistic overview of the current DLM landscape. We trace its evolution and relationship with other paradigms, such as autoregressive and masked language models, and cover both foundational principles and state-of-the-art models. Our work offers an up-to-date, comprehensive taxonomy and an in-depth analysis of current techniques, from pre-training strategies to advanced post-training methods. Another contribution of this survey is a thorough review of DLM inference strategies and optimizations, including improvements in decoding parallelism, caching mechanisms, and generation quality. We also highlight the latest approaches to multimodal extensions of DLMs and delineate their applications across various practical scenarios. Furthermore, our discussion addresses the limitations and challenges of DLMs, including efficiency, long-sequence handling, and infrastructure requirements, while outlining future research directions to sustain progress in this rapidly evolving field. Project GitHub is available at https://github.com/VILA-Lab/Awesome-DLMs.
@@ 1 Introduction
Recent advancements towards artificial general intelligence (AGI) have been largely driven by the emergence of autoregressive large language models (LLMs) [1, 2, 3, 4, 5, 6, 7] and diffusion models for image and video generation [8, 9, 10, 11, 12], These models exhibit remarkable capabilities in both understanding and generation across diverse modalities, achieving levels of performance that were previously unimaginable. The unprecedented scale of these models, reflected in massive parameter counts, vast datasets, substantial efforts in training, and significant computational demands during inference, has pushed AI to new heights, equipping these models with broad general knowledge and a deep understanding of language and the real world.
@@
To adapt diffusion for discrete language data, several key approaches have been proposed. In the early times, the development of DLMs was primarily driven by diffusion models’ success in continuous domains like image synthesis. Continuous DLMs map tokens into embeddings and perform denoising in continuous space, as in pioneer works Diffusion-LM [22] and SED [23]. Discrete DLMs, on the other hand, define the diffusion process directly in token space. Early efforts such as D3PM [24] introduced structured transition matrices with absorbing states, allowing token-level corruption and iterative denoising. Subsequent work like DiffusionBERT [25] integrated pre-trained masked language models (e.g., BERT) to enhance denoising quality, and proposed tailored noise schedules (e.g., the spindle schedule) to better align token corruption with token frequency. These early models demonstrated the feasibility of applying iterative denoising to non-autoregressive text generation, offering controllability and parallelism, though their performance still lagged behind strong autoregressive baselines. As core challenges in DLMs are gradually addressed and the paradigm matures, larger-scale DLMs have been developed. By initializing from autoregressive models, 7B-level models like Dream [26] and DiffuLLaMA [27] have shown that DLMs can be effectively adapted from existing models while achieving competitive performance. LLaDA-8B [28] further demonstrates the potential of training DLMs from scratch, achieving performance comparable to similarly sized LLaMA3-8B models. Multimodal DLMs, also known as diffusion multimodal large language models (dMLLMs), have also shown promise in modeling hybrid data such as text and images. Built upon open-source DLMs, models like LLaDA-V [29], Dimple [30], and MMaDA [31] integrate cross-modal reasoning and generation into the diffusion framework. Meanwhile, industry efforts have also shown growing interest in DLMs. The Mercury series [32] and Gemini Diffusion [33] report strong performance while achieving inference speeds of thousands of tokens per second. These developments highlight the growing practicality and commercial potential of DLMs. We provide a timeline of DLMs development in Fig. 1, ranging from representative models to recent advancements[34, 35, 36, 37, 38, 39, 40], followed by a visualization of DLM trends in Fig. 2.
@@
{forest}
for tree= grow=east, reversed, forked edges, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=gray, rounded corners, align=center, minimum width=6em, edge+=darkgray, line width=0.7mm, s sep=4pt, inner xsep=5pt, inner ysep=5pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=12em, font=, minimum height=3.5em, where level=2text width=13em, font=, minimum height=3.5em, where level=3 text width=40em, font=, align=left, fill=gray!10, text=black, inner xsep=6pt, inner ysep=6pt, , [ Taxonomy of Diffusion Language Models, ver, line width=0.7mm [ Paradigms (§ 2) , fill=c1!60, draw=c1, line width=0.5mm [ Continuous Space Models, fill=c1!60, draw=c1 [ Diffusion-LM [22], SED [23], LATENTOPS [49], Diffuseq [50], CDCD [51],
Difformer [52], LD4LG [53], GENIE [54], InfoDiffusion [55], EDDPMs [56],
SMOOTHIE [57], TESS [58], TESS 2 [59], LDEBM [60] ], ] [ Discrete Space Models, fill=c1!60, draw=c1, align=center [ D3PM [24], DiffusionBERT [25], LLaDA [28], RDMs [61],
MD4 [62], MDLM [63], Diffusion-LLM [64], Diffusion-NAT [65],
Plaid [66], SEDD [67], RADD [68], DFM [69],
DDPD [70], MGDM [71], Diffu-LLaMA [27], Dream-7B [26],
GIDD [72], LongLLaDA [73] ] ] [ Hybrid DLMs, fill=c1!60, draw=c1, align=center [ SSD-LM [74], AR-DIFFUSION [75], BD3-LM [76], CtrlDiff [77], SpecDiff [78]] ] ] [ Training Strategies (§ 3) , fill=c2!60, draw=c2, line width=0.5mm [ Pre-training, fill=c2!60, draw=c2 [ From scratch: LLaDA-8B [28]
Adapting from AR models: Dream [26], DiffuLLaMA [27]
Adapting from image diffusion models: D-DiT [79], Muddit [80] ] ] [ Post-training, fill=c2!60, draw=c2 [ DoT [81], DCoLT [82]
Policy Gradient: diffu-GRPO [42], UniGRPO [31], SEPO [83], Coupled-GRPO [84]
Preference Optimization: VRPO [85] ] ] ] [ Inference & Optimization (§ 4) , fill=c3!60, draw=c3, line width=0.5mm [ Parallel Decoding, fill=c3!60, draw=c3 [ Fast-dLLM [44], APD [45], SlowFast Sampling [86], SpecDiff [78], Dimple [30]] ] [ Unmasking/Remasking, fill=c3!60, draw=c3 [ LLaDA [28], Dream [26], Masked DLM [63], Fast-dLLM [44], ReMDM [46]] ] [ Guidance, fill=c3!60, draw=c3 [ A-CFG [87], Freecache [88], DINGO [89]] ] [ Efficiency Techniques, fill=c3!60, draw=c3 [ Key-Value Cache [44, 47, 48, 76];
Feature Cache [47, 88, 90, 91, 92, 93];
Step Distillation [94, 43, 95]; ] ] ] [ Multimodal & Applications (§ 5, § 7) , fill=c4!60, draw=c4, line width=0.5mm [ Multimodal DLMs, fill=c4!60, draw=c4 [ LLaDA-V [29], Dimple [30], MMaDA [31], D-DiT [79], LaViDa [96]
Fudoki [97], Muddit [80], UniDisc [98], ] ] [ Conventional NLP Tasks, fill=c4!60, draw=c4 [ ROIC-DM [99], DiffusionNER [100], IPAD [101], DiffusionABSA [102], DiffuSum [103]
TermDiffuSum [104], Diff-KPE [105], IPED [106], EdiText [107], DIFFUSEMP [108]
DiffuDetox [109], ParaGuide [110], PLANNER [111], DiffuCom [112], DiffusionDialog [113]
LDP [114], PoetryDiffusion [115], XDLM [116], DiffusionRet [117], DIFND [118]] ] [ Code Generation, fill=c4!60, draw=c4 [ DUS [119], DiffuCoder [84], DCoLT [82], Mercury Coder [32]] ] [ Computational Biology, fill=c4!60, draw=c4 [ Molecular : TransDLM [120], TGM-DLM [121]
Protein Design: MeMDLM [122], DPLM [123], CFP-GEN [124]
DRAKES [125], ForceGen [126], DSM [127], DPLM2 [128] ] ] ] ]
@@ 2.1.2 Autoregressive Language Models
The sequential generation formulation is both a strength and a limitation. On one hand, it aligns with text generation tasks and facilitates straightforward sampling, naturally suits for various applications. On the other hand, it imposes a fundamental bottleneck on inference speed, as token generation is inherently sequential and cannot be parallelized. This trade-off between generation quality and latency has become a central challenge in advancing AR models. Beyond the standard next-token prediction (NTP), recent research has explored multi-token prediction (MTP) [136, 16] to accelerate inference by generating multiple tokens per step. These efforts share conceptual similarities with parallel decoding strategies employed in DLMs.
@@ 2.1.3 Other Paradigms
$\mathcal{L}_{\text{PLM}}=\mathbb{E}_{\mathbf{z}\sim\mathcal{Z}_{T}}\left[-\sum_{t=1}^{N}\log P_{\theta}(x_{z_{t}}|\mathbf{x}_{\mathbf{z}_{<t}})\right]$ (5)
@@ 2.2 Continuous Diffusion Language Models
After training, generation proceeds by sampling from the learned reverse process, starting from noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. At each timestep $t=T,T-1,\ldots,1$, the model defines a conditional distribution $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ which aims to approximate the true reverse transition $q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$. Sampling iteratively from these learned conditionals produces progressively less noisy latent states until an estimate of the original data $\mathbf{x}_{0}$ is recovered. After generating a denoised embedding $\hat{\mathbf{x}}_{0}$, a rounding step maps it back to a discrete token. This is typically done by nearest-neighbor search in the embedding space or using a decoder head.
@@
Model Parameters Diffusion type Noise schedule Task Training data
D3PM [24] 70M Discrete Mutual information Language 65B tokens
Diffusion-LM [22] 100M & 300M Continuous Square-root Language
Diffuseq [50] 91M Continuous Square-root Language 565K sentence pairs
SSD-LM [74] 400M Continuous Cosine Language 123B tokens
DiffusionBERT [25] 110M Discrete Spindle Language 16B tokens
@@
LD4LG [53] 188M Continuous Cosine Language 5.2M sentence pairs
SeqDiffuSeq [140] 65M & 110M Continuous Adaptive Language 45B tokens
TESS [58] 125M & 355M Continuous Linear Language
MDLM [63] 110M Discrete Log-linear Language 622B tokens
DFM [69] 1.7B Discrete Linear & Cubic Language & Code 2.5T tokens
TESS-2 [59] 7B Continuous Log-linear Language 360B tokens
LLaDA [28] 1B & 8B Discrete Linear Language & Code 2.3T tokens
Mecury [32] Discrete Log-linear Code Trillions tokens
LLaDA-1.5 [85] 8B Discrete Linear Language 2.3T tokens
MMaDA [31] 8B Discrete Log-linear Language 900B image-text tokens
Dream [26] 7B Discrete Log-linear Language & Code 580B tokens
LLaDA-V [29] 8.4B Discrete Linear Multimodal 3M image-text samples
LaViDa [96] 8.4B Discrete Convex Multimodal 1.6M image-text samples
Dimple [30] 7B Discrete Log-linear Multimodal 0.8B tokens
LongLLaDA [73] 8B Discrete Log-linear Language & Code 2.3T tokens
DiffuCoder [84] 7B Discrete Log-linear Code 130B tokens
@@ 2.3 Discrete Diffusion Language Models
DiffusionBERT [25] combines a pre-trained BERT with a discrete diffusion process, leveraging its powerful denoising capabilities to learn the reverse process from a masked state. The model is further enhanced by a novel spindle noise schedule that considers token informativeness, achieving significant improvements in generation quality compared with previous DLMs. A different approach, Reparameterized Discrete diffusion Models (RDMs) [61], establishes an alternative formulation for the reverse process, which simplifies the training objective to a weighted cross-entropy loss. This enables more flexible and adaptive decoding strategies, leading to significant performance gains over previous discrete diffusion models. Similarly, MD4 [62] derives a simple weighted integral of cross-entropy losses as the continuous-time variational objective of masked diffusion models, providing a simple and generalized framework for training DLMs. Another analogous approach is MDLM [63], which introduces a simplified, Rao-Blackwellized objective that takes the form of a weighted average of masked language modeling losses. Diffusion-LLM [64] demonstrates the scalability of DLMs by adapting pre-trained masked language models to diffusion paradigm and further task-specific finetuning and instruction finetuning, unlocking their versatility in solving general language tasks. Diffusion-NAT [65] unifies a discrete diffusion model with a PLM by reformulating the denoising process as a non-autoregressive masked token recovery task, allowing BART to act as an effective denoiser. Plaid [66] is the first diffusion language model trained to maximize data likelihood, demonstrating through scaling laws that it can outperform autoregressive models like GPT-2 on standard benchmarks. To improve the training objective, SEDD [67] introduces a score entropy loss to directly learn the ratios of the data distribution, which serves as a discrete extension of score matching. Reparameterized Absorbing Discrete Diffusion (RADD) [68] reveals that the concrete score in absorbing diffusion can be expressed as a time-independent conditional probability of the clean data, multiplied by an analytic, time-dependent scalar. It also formally unifies the training objectives of absorbing discrete diffusion and any-order autoregressive models. Discrete Flow Matching (DFM) [69] introduces a novel generative paradigm for discrete data that is analogous to continuous Flow Matching. The method learns a generating probability velocity to transform samples along a general family of probability paths from a source to a target distribution. By scaling the model architecture, DFM significantly closes the performance gap with autoregressive models on various benchmarks. DDPD [70] presents a framework that decouples the generation process into two specialized models: a planner and a denoiser. At each step, the planner identifies the most corrupted token positions needing refinement, after which the denoiser predicts their values. To improve performance on complex reasoning tasks, MGDM [71] is introduced to address the problem of subgoal imbalance. This approach enhances discrete diffusion by prioritizing more difficult subgoals during the learning process through a token-level reweighting mechanism. To address the challenge of scaling, a continual pre-training approach [27] is proposed to adapt existing autoregressive models, such as LLaMA, into diffusion language models. The resulting models, named DiffuGPT and DiffuLLaMA, are competitive with their AR counterparts while gaining diffusion-native capabilities like flexible infilling. Build on this observation, Dream-7B [26] is initialized from Qwen2.5 7B [142] and further trained with 580B tokens, largely outperforming existing DLMs and matching the performance of top-tier AR models. GIDD [72] is introduced to overcome the limitation that masked diffusion models cannot revise generated tokens. This framework generalizes the noising process by combining masking with uniform noise, which unlocks the model’s ability to self-correct mistakes and improves sample quality. Recently, to address long-context capabilities, LongLLaDA [73] provides the first systematic analysis of DLMs in this domain. It reveals that DLMs can maintain stable perplexity during direct context extrapolation and have better retrieval capabilities. LongLLaDA also introduces a training-free NTK-based RoPE extrapolation method, which significantly improves the extrapolation performance of DLMs, validating that established extrapolation scaling laws remain effective for DLMs.

### v2

@@ Abstract
Diffusion Language Models (DLMs) are rapidly emerging as a powerful and promising alternative to the dominant autoregressive (AR) paradigm. By generating tokens in parallel through an iterative denoising process, DLMs possess inherent advantages in reducing inference latency and capturing bidirectional context, thereby enabling fine-grained control over the generation process. While achieving several-fold speedups, recent advancements have allowed DLMs to show performance comparable to their autoregressive counterparts, making them a compelling choice for various natural language processing tasks. Despite their growing prevalence, DLMs present challenges and opportunities that warrant further exploration, requiring a detailed and systematic understanding of their principles, techniques, and limitations. In this survey, we provide a holistic overview of the current DLM landscape. We trace its evolution and relationship with other paradigms, such as autoregressive and masked language models, and cover both foundational principles and state-of-the-art models. Our work offers an up-to-date, comprehensive taxonomy and an in-depth analysis of current techniques, from pre-training strategies to advanced post-training methods. Another contribution of this survey is a thorough review of DLM inference strategies and optimizations, including improvements in decoding parallelism, caching mechanisms, and generation quality. We also highlight the latest approaches to multimodal extensions of DLMs and delineate their applications across various practical scenarios. Furthermore, our discussion addresses the limitations and challenges of DLMs, including efficiency, long-sequence handling, and infrastructure requirements, while outlining future research directions to sustain progress in this rapidly evolving field. Project GitHub is available at https://github.com/VILA-Lab/Awesome-DLMs.
@@ 1 Introduction
Recent advancements toward artificial general intelligence (AGI) have been largely driven by the emergence of autoregressive large language models (LLMs) [1, 2, 3, 4, 5, 6, 7] and diffusion models for image and video generation [8, 9, 10, 11, 12]. These models exhibit remarkable capabilities in both understanding and generation across diverse modalities, achieving levels of performance that were previously unimaginable. The unprecedented scale of these models, reflected in massive parameter counts, vast datasets, substantial efforts in training, and significant computational demands during inference, has pushed AI to new heights, equipping these models with broad general knowledge and a deep understanding of language and the real world.
@@
To adapt diffusion for discrete language data, several key approaches have been proposed. In the early times, the development of DLMs was primarily driven by diffusion models’ success in continuous domains like image synthesis. Continuous DLMs map tokens into embeddings and perform denoising in continuous space, as in pioneering works Diffusion-LM [22] and SED [23]. Discrete DLMs, on the other hand, define the diffusion process directly in token space. Early efforts such as D3PM [24] introduced structured transition matrices with absorbing states, allowing token-level corruption and iterative denoising. Subsequent work like DiffusionBERT [25] integrated pre-trained masked language models (e.g., BERT) to enhance denoising quality, and proposed tailored noise schedules (e.g., the spindle schedule) to better align token corruption with token frequency. These early models demonstrated the feasibility of applying iterative denoising to non-autoregressive text generation, offering controllability and parallelism, though their performance still lagged behind strong autoregressive baselines. As core challenges in DLMs are gradually addressed and the paradigm matures, larger-scale DLMs have been developed. By initializing from autoregressive models, 7B-level models like Dream [26] and DiffuLLaMA [27] have shown that DLMs can be effectively adapted from existing models while achieving competitive performance. LLaDA-8B [28] further demonstrates the potential of training DLMs from scratch, achieving performance comparable to similarly sized LLaMA3-8B models. Multimodal DLMs, also known as diffusion multimodal large language models (dMLLMs), have also shown promise in modeling hybrid data such as text and images. Built upon open-source DLMs, models like LLaDA-V [29], Dimple [30], and MMaDA [31] integrate cross-modal reasoning and generation into the diffusion framework. Meanwhile, industry efforts have also shown growing interest in DLMs. The Mercury series [32], Gemini Diffusion [33], and Seed Diffusion [34] report strong performance while achieving inference speeds of thousands of tokens per second. These developments highlight the growing practicality and commercial potential of DLMs. We provide a timeline of DLMs development in Fig. 1, ranging from representative models to recent advancements[35, 36, 37, 38, 39, 40, 41], followed by a visualization of DLM trends in Fig. 2.
@@
Taxonomy of Diffusion Language Models
Paradigms (§ 2) Continuous Space Models Diffusion-LM [22], SED [23], LATENTOPS [50], Diffuseq [51], CDCD [52], Difformer [53], LD4LG [54], GENIE [55], InfoDiffusion [56], EDDPMs [57], SMOOTHIE [58], TESS [59], TESS 2 [60], LDEBM [61] Discrete Space Models D3PM [24], DiffusionBERT [25], LLaDA [28], RDMs [62], MD4 [63], MDLM [64], Diffusion-LLM [65], Diffusion-NAT [66], Plaid [67], SEDD [68], RADD [69], DFM [70], DDPD [71], MGDM [72], Diffu-LLaMA [27], Dream-7B [26], GIDD [73], LongLLaDA [74], Seed Diffusion [34], LLaDA-MoE [75] Hybrid DLMs SSD-LM [76], AR-DIFFUSION [77], BD3-LM [78], CtrlDiff [79], SpecDiff [80], SDAR [81], TiDAR [82], SDLM [83] Training Strategies (§ 3) Pre-training From scratch: LLaDA-8B [28] Adapting from AR models: Dream [26], DiffuLLaMA [27] Adapting from image diffusion models: D-DiT [84], Muddit [85] Post-training DoT [86], DCoLT [87] Policy Gradient: diffu-GRPO [43], UniGRPO [31], SEPO [88], Coupled-GRPO [89], wd1 [90], IGPO [91], SPG [92], SAPO [93], BGPO [94] Preference Optimization: VRPO [95] Inference & Optimization (§ 4) Parallel Decoding Fast-dLLM [45], APD [46], SlowFast Sampling [96], SpecDiff [80], Dimple [30] Learn2PD [97], dParallel [98] Unmasking/Remasking LLaDA [28], Dream [26], Masked DLM [64], Fast-dLLM [45], ReMDM [47] Guidance A-CFG [99], Freecache [100], DINGO [101] Efficiency Techniques Key-Value Cache [45, 48, 49, 78, 102, 103]; Feature Cache [48, 100, 104, 105, 106, 107]; Step Distillation [108, 44, 109]; Multimodal & Applications (§ 5, § 7) Multimodal DLMs LLaDA-V [29], Dimple [30], MMaDA [31], D-DiT [84], LaViDa [110], Fudoki [111] Muddit [85], UniDisc [112], Lumina-DiMOO [113], LaViDa-O [114], MMaDA-Parallel [115] Conventional NLP Tasks ROIC-DM [116], DiffusionNER [117], IPAD [118], DiffusionABSA [119], DiffuSum [120] TermDiffuSum [121], Diff-KPE [122], IPED [123], EdiText [124], DIFFUSEMP [125] DiffuDetox [126], ParaGuide [127], PLANNER [128], DiffuCom [129], DiffusionDialog [130] LDP [131], PoetryDiffusion [132], XDLM [133], DiffusionRet [134], DIFND [135] Code Generation DUS [136], DiffuCoder [89], DCoLT [87], Mercury Coder [32] Computational Biology Molecular : TransDLM [137], TGM-DLM [138] Protein Design: MeMDLM [139], DPLM [140], CFP-GEN [141] DRAKES [142], ForceGen [143], DSM [144], DPLM2 [145] Robotics LLaDA-VLA [146], dVLA [147], UD-VLA [148]
@@ 2.1.2 Autoregressive Language Models
The sequential generation formulation is both a strength and a limitation. On one hand, it aligns with text generation tasks and facilitates straightforward sampling, naturally suits for various applications. On the other hand, it imposes a fundamental bottleneck on inference speed, as token generation is inherently sequential and cannot be parallelized. This trade-off between generation quality and latency has become a central challenge in advancing AR models. Beyond the standard next-token prediction (NTP), recent research has explored multi-token prediction (MTP) [156, 16] to accelerate inference by generating multiple tokens per step. These efforts share conceptual similarities with parallel decoding strategies employed in DLMs, while some other works are directly inspired by diffusion process to align LLMs [157].
@@ 2.1.3 Other Paradigms
$\mathcal{L}_{\text{PLM}}=\mathbb{E}_{\mathbf{z}\sim\mathcal{Z}_{T}}\left[-\sum_{t=1}^{T}\log P_{\theta}(x_{z_{t}}|\mathbf{x}_{\mathbf{z}_{<t}})\right]$ (5)
@@ 2.2 Continuous Diffusion Language Models
After training, generation proceeds by sampling from the learned reverse process, starting from noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. At each timestep $t=T,T-1,\ldots,1$, the model defines a conditional distribution $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ which aims to approximate the true reverse transition $q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$. Sampling iteratively from these learned conditionals produces progressively less noisy latent states until an estimate of the original data $\mathbf{x}_{0}$ is recovered. After generating a denoised embedding $\hat{\mathbf{x}}_{0}$, a rounding step maps it back to a discrete token. This is typically done by nearest-neighbor searching in the embedding space, using a decoder head, or thresholding techniques [163].
@@
Model Parameters Diffusion type Task Training data
D3PM [24] 70M Discrete Language 65B tokens
Diffusion-LM [22] 100M & 300M Continuous Language
Diffuseq [51] 91M Continuous Language 565K sentence pairs
SSD-LM [76] 400M Continuous Language 123B tokens
DiffusionBERT [25] 110M Discrete Language 16B tokens
@@
LD4LG [54] 188M Continuous Language 5.2M sentence pairs
SeqDiffuSeq [161] 65M & 110M Continuous Language 45B tokens
TESS [59] 125M & 355M Continuous Language
MDLM [64] 110M Discrete Language 622B tokens
DFM [70] 1.7B Discrete Language & Code 2.5T tokens
TESS-2 [60] 7B Continuous Language 360B tokens
LLaDA [28] 1B & 8B Discrete Language & Code 2.3T tokens
Mercury [32] Discrete Code Trillions of tokens
LLaDA-1.5 [95] 8B Discrete Language 2.3T tokens
MMaDA [31] 8B Discrete Multimodal Unified 900B image-text tokens
Dream [26] 7B Discrete Language & Code 580B tokens
LLaDA-V [29] 8.4B Discrete Multimodal 3M image-text samples
LaViDa [110] 8.4B Discrete Multimodal 1.6M image-text samples
Dimple [30] 7B Discrete Multimodal 0.8B tokens
LongLLaDA [74] 8B Discrete Language & Code 2.3T tokens
DiffuCoder [89] 7B Discrete Code 130B tokens
LaViDa-O [114] 10.4B Discrete Multimodal Unified >> 200M image-text pairs
Lumina-DiMOO [113] 8B Discrete Multimodal Unified >> 110M image-text pairs
LLaDA-MoE [75] 7B (1.4B activated) Discrete Language 20T tokens
SDAR [81] 1.7B-30B series Discrete Language 54B tokens
TiDAR [82] 1.5B & 8B Discrete Language 150B tokens
SDLM [83] 3B & 32B Discrete Language 2.3B tokens
@@ 2.3 Discrete Diffusion Language Models
DiffusionBERT [25] combines a pre-trained BERT with a discrete diffusion process, leveraging its powerful denoising capabilities to learn the reverse process from a masked state. The model is further enhanced by a novel spindle noise schedule that considers token informativeness, achieving significant improvements in generation quality compared with previous DLMs. A different approach, Reparameterized Discrete diffusion Models (RDMs) [62], establishes an alternative formulation for the reverse process, which simplifies the training objective to a weighted cross-entropy loss. This enables more flexible and adaptive decoding strategies, leading to significant performance gains over previous discrete diffusion models. Similarly, MD4 [63] derives a simple weighted integral of cross-entropy losses as the continuous-time variational objective of masked diffusion models, providing a simple and generalized framework for training DLMs. Another analogous approach is MDLM [64], which introduces a simplified, Rao-Blackwellized objective that takes the form of a weighted average of masked language modeling losses. Diffusion-LLM [65] demonstrates the scalability of DLMs by adapting pre-trained masked language models to diffusion paradigm and further task-specific finetuning and instruction finetuning, unlocking their versatility in solving general language tasks. Diffusion-NAT [66] unifies a discrete diffusion model with a PLM by reformulating the denoising process as a non-autoregressive masked token recovery task, allowing BART to act as an effective denoiser. Plaid [67] is the first diffusion language model trained to maximize data likelihood, demonstrating through scaling laws that it can outperform autoregressive models like GPT-2 on standard benchmarks. To improve the training objective, SEDD [68] introduces a score entropy loss to directly learn the ratios of the data distribution, which serves as a discrete extension of score matching. Reparameterized Absorbing Discrete Diffusion (RADD) [69] reveals that the concrete score in absorbing diffusion can be expressed as a time-independent conditional probability of the clean data, multiplied by an analytic, time-dependent scalar. It also formally unifies the training objectives of absorbing discrete diffusion and any-order autoregressive models. Discrete Flow Matching (DFM) [70] introduces a novel generative paradigm for discrete data that is analogous to continuous Flow Matching. The method learns a generating probability velocity to transform samples along a general family of probability paths from a source to a target distribution. By scaling the model architecture, DFM significantly closes the performance gap with autoregressive models on various benchmarks. DDPD [71] presents a framework that decouples the generation process into two specialized models: a planner and a denoiser. At each step, the planner identifies the most corrupted token positions needing refinement, after which the denoiser predicts their values. To improve performance on complex reasoning tasks, MGDM [72] is introduced to address the problem of subgoal imbalance. This approach enhances discrete diffusion by prioritizing more difficult subgoals during the learning process through a token-level reweighting mechanism. To address the challenge of scaling, a continual pre-training approach [27] is proposed to adapt existing autoregressive models, such as LLaMA, into diffusion language models. The resulting models, named DiffuGPT and DiffuLLaMA, are competitive with their AR counterparts while gaining diffusion-native capabilities like flexible infilling. Build on this observation, Dream-7B [26] is initialized from Qwen2.5 7B [164] and further trained with 580B tokens, largely outperforming existing DLMs and matching the performance of top-tier AR models. GIDD [73] is introduced to overcome the limitation that masked diffusion models cannot revise generated tokens. This framework generalizes the noising process by combining masking with uniform noise, which unlocks the model’s ability to self-correct mistakes and improves sample quality. Recently, to address long-context capabilities, LongLLaDA [74] provides the first systematic analysis of DLMs in this domain. It reveals that DLMs can maintain stable perplexity during direct context extrapolation and have better retrieval capabilities. LongLLaDA also introduces a training-free NTK-based RoPE extrapolation method, which significantly improves the extrapolation performance of DLMs, validating that established extrapolation scaling laws remain effective for DLMs. UltraLLaDA [165] extends this line of work by introducing a diffusion-aware NTK RoPE scaling and lightweight long-context post-training, enabling diffusion LLMs to reach 128K context windows and achieving substantially better retrieval and perplexity performance than training-free extrapolation methods. LLaDA-MoE [75] is the first work to integrate a sparse Mixture-of-Experts (MoE) architecture into diffusion language models, training a new MoE-based DLM from scratch on 20T tokens. Despite activating only about 1.4B parameters during inference, it surpasses larger dense diffusion models and achieves performance comparable to Qwen2.5-3B-Instruct across knowledge, coding, and reasoning benchmarks.